OpenStudy (anonymous):

Find the difference quotient and simplify your answer. g(x)=1/x^2 , g(x) - g(3) / x-3 , x does not equal 3

OpenStudy (anonymous):

so far i have \[\frac{ \frac{ 1 }{ x^2 } - \frac{ 1 }{ 9 } }{ x-3}\] but I don't know what to do now.

OpenStudy (anonymous):

well i believe you can expand your equation out\[\frac{ 1 }{ x ^{2} }-\frac{ 1 }{ 9 } /x-3\] this is the the same thing as

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{2}} - \frac{ 1 }{ 9 } * x-3\] hopefully that helps and you can do the rest. If not ill help

OpenStudy (anonymous):

I forgot how to multiply those together :/ sorry, could you tell me how?

OpenStudy (anonymous):

First.. Dont multiply yet.. find common denominators for the two fractions then subtract across.. then youll have one single fraction multiplied by the x-3.. then just multiply the numerator by the x-3

OpenStudy (anonymous):

\[\frac{ 9-x^2 }{ x-3 }\] ? /:

OpenStudy (anonymous):

Ok good (im not really checking your work I can though if you want..) now factor the numerator) and simplify

OpenStudy (anonymous):

I got: \[\frac{ -(x+3) }{ 9x^2 }\] & I'd really appreciate it if you could check my work because i didn't get the answer that matched what was in the book /: the answer should be \[\frac{ x+3 }{ 9x^2 }\]

OpenStudy (anonymous):

OK in the beginning is the question asking for everything to be divided by x-3 or only by g(3)?

OpenStudy (anonymous):

Everything to be divided.

OpenStudy (anonymous):

This is the problem (i rewrote it so it would be easier to see if that helps) : \[g(x)= \frac{ 1 }{ x^2 }, \frac{ g(x)-g(3) }{ x-3 }\]

OpenStudy (anonymous):

I do not get what the book gets for some reason

OpenStudy (anonymous):

Yeah, I did the same thing you did. /: It's just that negative that's messing everything up. Blahhhh. Well, I guess I'll just ask my teacher on Monday. Thanks! (:

OpenStudy (anonymous):

Youre very welcome! sorry i couldnt completely help you out